### v2

# A Buddy for Betelgeuse: Binarity as the Origin of the Long Secondary Period in Ā Orionis

Abstract
We predict the existence of Ā OriĀ B, a low-mass companion orbiting Betelgeuse. This is motivated by the presence of a 2170-day Long Secondary Period (LSP) in Betelgeuseās lightcurve, a periodicity times longer than the starās 416 day fundamental radial pulsation mode. While binarity is currently the leading hypothesis for LSPs in general, the LSP and the radial velocity variation observed in Betelgeuse, taken together, necessitate a revision of the prevailing physical picture. Specifically, the lightcurve-RV phase difference requires a companion to be behind Betelgeuse at the LSP luminosity minimum, 180 degrees out of phase with the system orientation associated with occultation. We demonstrate the consistency of this model with available observational constraints and identify tensions in all other proposed LSP hypotheses. Within this framework, we calculate a mass for Ā OriĀ B of and an orbital separation of , or times the radius of Betelgeuse. We then describe the features of the companion as constrained by the fundamental parameters of Betelgeuse and its orbital system, and discuss what would be required to confirm the companionās existence observationally.
1 Introduction
Interest in Betelgeuse, more formally known as Orionis, has blossomed over the past five years due in large part to its anomalously severe dimming observed between the end of 2019 and beginning of 2020 (Guinan etĀ al., 2019; Harper etĀ al., 2020; Dupree etĀ al., 2020). Now widely understood to have been caused by a dust cloud (MontargĆØs etĀ al., 2021), the āGreat Dimmingā nonetheless inspired many groundbreaking investigations into our nearest supergiant. These studies have led to revisions in our understanding of Betelgeuseās behavior and fundamental parameters, as well as opened new lines of inquiry into some of Betelgeuseās less well-understood properties, such as its apparent 5 km/s surface rotation rate (Ma etĀ al., 2024), the characterization of its variability (Jadlovský etĀ al., 2024), and the classification of its pulsation modes (Joyce etĀ al., 2020; Saio etĀ al., 2023; MacLeod etĀ al., 2023).
A distinctive variability cycle observed in Betelgeuse is a 2100 day period in the lightcurve, (see, e.g. Kiss etĀ al., 2006; Joyce etĀ al., 2020; Granzer etĀ al., 2022; MacLeod etĀ al., 2023), often labelled a Long Secondary Period (LSP). LSPs are seen in many cool, luminous stars ranging from Red Giant Branch stars (RGB; e.g.Ā Wood etĀ al. 1999a), Asymptotic Giant Branch stars (AGB; e.g.Ā Zijlstra & Bedding 2002; Uttenthaler etĀ al. 2015), post-AGB stars (Van Winckel etĀ al., 1999; Bódi etĀ al., 2016), Yellow Hypergiants (YHGs; e.g. Percy & Kim 2014), and Red Supergiant (RSG; e.g. Kiss etĀ al. 2006) stars, in all cases showing generally steady periods ranging from hundreds to thousands of days (e.g.Ā Takayama 2023). Long Secondary Periods are an as-yet unexplained class of stellar variability, called āsecondaryā because LSPs are typically a few to ten times slower than their host starsā radial fundamental pulsation modes (FMs).
Outstanding among the questions raised by renewed fascination with and deeper investigation into Betelgeuse is whether Betelgeuseās prominent, 2100-day periodicity is an LSP or the FM. The answer to this question carries significant implications for Betelgeuseās current evolutionary phase, which in turn implies the timeline for Betelgeuseās death in a supernova explosion (Wheeler etĀ al., 2017; Nance etĀ al., 2018; Chatzopoulos etĀ al., 2020; Saio etĀ al., 2023). If the 2100-day periodicity is the fundamental mode, it implies a large and observationally contentious radius for Betelgeuse. Further, it would place Betelgeuseās current evolutionary stage beyond the onset of core carbon burning, suggesting that a supernova explosion is imminent within the next several dozen to several hundred years (Saio etĀ al., 2023). If, on the other hand, the 2100-d periodicity is an LSP, Betelgeuse is comfortably within in its core helium burning phase and not due for an explosion for hundreds of thousands of years (Joyce etĀ al., 2020). As we discuss below, most studies directly or indirectly support classifying the 420-day periodicity as the fundamental mode. As such, the cause of Betelgeuseās 2100-day periodicity remains elusive.
There have been several mechanisms proposed to explain LSPs in general, ranging from giant cell convection, mode interactions, non-radial pulsations, and binarity, as well as more exotic physics, such as dust modulation beyond the stellar surface (many of these are discussed in Wood etĀ al. 2004a). We explore the relative merits of each of these and more hypotheses proposed in the literature individually in the context of modern observational data (see Section 4). We ultimately identify binarity as the most likely among them.
The idea that binarity could explain LSPs dates back to the late 1990s and early 2000s (Wood etĀ al., 1999a, 2004b) and has been reinforced since by the sequences formed by variable stars on period-luminosity diagrams (SoszyÅski, 2007) and in RV data (Hinkle etĀ al., 2002; Nicholls etĀ al., 2009). While such early versions of the binarity hypothesis were most concerned with whether a close companion could induce low-frequency modes on the primary star, the current leading theory is that the timescale of the LSP period is set by the orbital time of a low-mass companion, and the mechanism of dimming and brightening involves the formation and removal of dust along the line of sight in phase with the companionās orbit (SoszyÅski etĀ al., 2021). Additionally, it has been proposed before that Betelgeuse is a multiple system (Karovska etĀ al., 1986) from speckle-imaging measurements, inferred to be two high-eccentricity companions including one at a separation of 0.06ā with a period of 2.1 years, which has not been confirmed in follow-up work with higher-resolution instruments (e.g. Wilson etĀ al., 1992; Kervella etĀ al., 2009; MontargĆØs etĀ al., 2016).
In this paper, we explore the implications and requirements of the binarity-as-LSP model for Ā Orionis. In Section 2, we discuss LSPs in cool, evolved stars in the context of the period-luminosity sequences formed by classical variable stars and consider where Betelgeuse fits in to this picture. In Section 3, we compare the brightness and radial velocity variations observed in Betelgeuse and a significant subset of Gaia LSP stars and discuss the physical implications of the LC-RV phase offsets. In Section 4, we overview all of the scenarios proposed as an explanation for Betelgeuseās LSP, demonstrating critical flaws in all but one case: Betelgeuse has a companion that interacts with the starās dusty circumstellar environment. In Section 5, we derive the properties this companion must have based on observational constraints (5.1) and discuss the conditions under which it should be observable (SectionĀ 5.2). We summarize and present our conclusions in Section 6.
2 LSPs in Cool Evolved Stars
LSPs, sometimes referred to as in older literature, are a common but not ubiquitous phenomenon in cool evolved stars. They have been observed in of long-period variables (LPVs; see, e.g. SoszyÅski etĀ al., 2021; Pawlak etĀ al., 2024) and have been detected both in the Milky Way and in the Magellanic Clouds (see, e.g. Payne-Gaposchkin, 1954; Wood etĀ al., 1999a; Kiss etĀ al., 2006; SoszyÅski etĀ al., 2009, 2011, 2013). These stars have large, radially extended convective envelopes, and with increasing luminosity, the convection in their envelopes can become near- or trans-sonic, with large-scale convective plumes spanning large fractions of the stellar radius (see, e.g., review by Chiavassa etĀ al., 2024). Stars hosting Long Secondary Periods by definition show other pulsation modes, which are identified as low-order radial pressure modes (e.g.ā fundamental and first-overtone, or FM and O1).
Large sky surveys targeting the Magellanic Clouds, such as the MACHO and OGLE surveys (Alcock etĀ al., 1997; Udalski etĀ al., 1992; Udalski, 2003), revealed that variability in red giant stars fall onto multiple, nearly parallel period-luminosity (P-L) relations, which were originally labeled from A to D (Wood etĀ al., 1999a). Sequences A, B and C were attributed to pulsation modes, whereas a fifth sequence, E, was attributed to eclipsing binaries (Derekas etĀ al., 2006). Sequence D included the Long Secondary Periods, which were found to be too long to be radial pulsation modes. We reproduce the sequences for the LMC from the OGLE-III survey in FigureĀ 1, which includes the later-identified Aā and Cā sequences (SoszyÅski etĀ al., 2009). Here we plot the absolute K magnitudes of the stars, after cross-matching the OGLE catalog with the IRSF Magellanic Clouds Point Source Catalog (Kato etĀ al., 2007) using a distance modulus of Ā mag (Wielgórski etĀ al., 2022).
Classical pulsating stars have been known to form P-L relations thanks to the pioneering work of Leavitt & Pickering (1912). These extend from classical Cepheids to other classes, such as Type II Cepheids, RRĀ Lyrae and Ā Scuti stars, but also to red giants, semiregular and Mira stars (Wyrzykowski etĀ al., 2012). Many authors, e.g., Yu etĀ al. (2020) and Trabucchi etĀ al. (2017), have studied how the red giant sequences relate to specific modes. The latter work concluded that sequence Aā, A and B correspond to the third, second and first radial overtones, respectively, while Cā and C separate less clearly due to the combined presence of RGB and AGB stars. Cā includes both overtone and fundamental-mode stars, whereas C is largely composed of Mira stars, which are AGB stars pulsating in the fundamental mode (e.g.Ā MolnĆ”r etĀ al. 2019; Joyce etĀ al. 2024).
LSPs in sequence D also follow a period-luminosity relationship that is distinct from the P-L relations discussed so far. Sequence D was discovered by Wood etĀ al. (1999a) using MACHO data. Binarity was proposed as a possible mechanism for the LSP by various authors (Hinkle etĀ al., 2002; Olivier & Wood, 2003; Wood etĀ al., 2004b). This was reaffirmed when Derekas etĀ al. (2006) showed that sequence D is an extension of sequence E (eclipsing binaries).
These sequences can also be extended further into the RSG regime (also shown in FigureĀ 1). Kiss etĀ al. (2006) found that Galactic supergiants form sequences that correspond to the O1, FM and LSP periodicities and extend from sequences B, C, and D, respectively. We recalculated the and absolute brightnesses of this RSG sample, using Gaia DR3 distances from the catalog of Bailer-Jones etĀ al. (2021) and plot these as blue filled circles in FigureĀ 1. These distances were sufficient in most cases, but a few very bright stars required other sources for parallax (van Leeuwen, 2007; Reid etĀ al., 2019). For Betelgeuse, we used the distance calculated by Joyce etĀ al. (2020), shown as red filled circles. Here we show Wesenheit indices as a proxy for luminosity based on the (J-K) colors, using the equation: , because that limits the effects of interstellar reddening and helps to separate the sequences better (Lebzelter etĀ al., 2019).
Compared to the original plot by Kiss etĀ al. (2006), the O1 and FM sequences reproduced in FigureĀ 1 appear to be considerably tighter. However, mode assignment based on PL or PW relations can be complicated as the constituent stars may belong to very different evolutionary stages, luminosity regimes, and variability classes that follow different relations. To illustrate this, we include three different PW relations with blue dashed lines, computed by SoszyÅski etĀ al. (2007), who identified further sequences and substructures in the red giant sample. The PW relation running through sequence C belongs to the Mira stars, whereas a1 and b1 belong to the lower-luminosity AGB and RGB stars, respectively. The FM supergiants lie close to the extrapolations of these relations. We discuss modal assignment in more detail and based on further arguments in Sect.Ā 4.1.
The RSG LSP periods in FigureĀ 1 show larger scatter and have larger uncertainties, but appear to form a coherent extension of sequence D. The periods identified in Betelgeuse as the O1, FM and LSP align with the RSG sequences. In the case of radial pressure modes, in which the pulsation period is determined by the stellar radius and the average sound speed, a P-L relation is expected, as the period scales with average density, thereby radius and mass, and thereby luminosity. In this case, the longest possible radial pressure mode is the FM. The apparent existence of a P-L relation in the LSP, on the other hand, is puzzling prima facie. Gravity modes can yield periods below the FM, but no such period-luminosity relation exists for g-modes.
Unless the LSP phenomenon is simply a ubiquitous off-by-one error in the classification of the FM (see SectionĀ 4.1), LSPs appear impossible to explain with conventional pulsation theory when considering only material within the stellar photosphere as the āstar.ā Whether the LSP is a pulsation or something else entirely, its period must be set by something other than the acoustic timescale of the star itself. As such, the LSP must operate on timescales longer than the dynamical timescale at the stellar surface , which is equivalent to the timescale of a Keplerian orbit at a distance equal to the stellar radius.
It is reasonable to ask why something without an intrinsic source of variability would nonetheless display a P-L relation. For this we can look outside the surface of the star, and make a dynamical argument. Generically, for a virial fluid element or body outside the stellar radius, a period is set by the Keplerian orbital timescale at that location. Although a generic orbit is not sensitive to the stellar luminosity, any orbital architecture knows about the stellar mass and radius, as , where is the distance from the center of the star to the external location setting the long secondary period timescale, i.e. the orbital separation.
If scales with , then we can estimate
where the last proportionality makes use of the approximations that (1) for gas-pressure-dominated stars and (2) . This is appropriate since the surface temperature range for RSGs, RGBs, and AGBs does not vary substantially relative to . Notably, the zero-point of this relation would depend on the exact relationship between and , which introduces scatter in the scaling relation. This crude scaling argument nonetheless illustrates that any period sensitive to a starās radius and mass, such as one set by dynamics outside the stellar surface, is in turn sensitive to the starās luminosity.
A notable feature of observed LSPs is that they appear more prominently in bluer bands, though the effective temperatures of their host stars are not thought to change dramatically (Takayama etĀ al., 2015; Gaia Collaboration etĀ al., 2023a). This suggests that LSPs are likely related to the formation, destruction, modulation, or attenuation of dust around stars. However, dust alone cannot explain the LSP phenomenon; the presence of a dust-modulating mechanism is also required. When searching in visual passbands, SoszyÅski etĀ al. (2021) found that only the attenuation by the cloud was visible. However, they identified secondary eclipses at longer wavelengths from WISE photometry, which is capable of detecting infrared radiation from the dust cloud directly. SoszyÅski etĀ al. (2021) thus conclude that LSPs arise when a small companion orbits the star together with an extended dust cloud.
Despite what we know about LSPs observationally and what we can infer about them theoretically, they remain to date the only form of steady-state stellar variability without a strong consensus on the underlying physics.
3 Characterizing the Lightcurve-RV Phase Offset
An important diagnostic for the nature of pulsational versus orbital behavior is the presence or lack of radial velocity (RV) modulations and the phase relative to brightness fluctuations.
Figure 2 shows the recent lightcurve (LC) and RV variations for Betelgeuse side by side. The photometric data include V-band observations collected by the American Association of Variable Star Observers (AAVSO) as well as the processed and rectified photometry of the star from the SMEI instrument (Jackson etĀ al., 2004; Hick etĀ al., 2007), originally published by Joyce etĀ al. (2020). RV observations were collected for more than a decade with the STELLA robotic telescope at Tenerife (Granzer etĀ al., 2022). From the latter, an LSP period of Ā d and a peak-to-peak amplitude of km/s were determined.
We fitted both variations with the  d periodicity, from which we calculate the phase offset (Granzer et al., 2022; Jadlovský et al., 2023). This period value is consistent with other observations (Kiss et al., 2006; Joyce et al., 2020).
| ID | P (d) | (mag) | (km/s) | e | |
| LSP | 2170 | 0.084(4) | 2.62(2) | 0.027 | |
| FM | 380 | 0.088(4) | 1.45(2) | 0.058 | |
| O1 | 216 | 0.075(4) | 1.72(2) | 0.059 |
To determine the phase offset between the LC and RV for Betelgeuse, we follow the phase lag method described by Szabó et al. (2007), where we calculate from the phases of the fundamental component of the following Fourier fit to the data:
where is either the average brightness or the average (also called ) velocity, is the common frequency between the two data sets, and and are the Fourier amplitude and phase components, respectively. For simplicity, we limit the fit to . However, we calculate the phase offset in Betelgeuse not just for the LSP but for the FM and O1 modes as well, based on the -band and SMEI curves from Joyce et al. (2020), plus newer -band observations collected by the AAVSO, truncated to the interval of the RV observations. We list our results for Betelgeuse in Table 1. The light curve and pulsation periods of Betelgeuse are known to vary, and they went through considerable changes after the Great Dimming again, with O1 becoming the dominant pulsation mode instead of the FM (Dupree et al., 2022; Jadlovský et al., 2024). With the incorporation of new observations, the average periods of both modes shifted slightly relative to the values found by Joyce et al. (2020): the O1 became longer, whereas the FM became shorter, matching the period found by Kiss et al. (2006).
Figure 3 shows the phase offsets between Betelgeuseās lightcurve and RV curve in the context of similar stars measured with Gaia. We compare Betelgeuse to a limited sample of LPV-candidate Gaia stars from Gaia Collaboration etĀ al. (2023a) for which we could estimate the phase offsets. We also show the pulsational phase lags for classical Cepheids for comparison, including high-precision data for 87 stars from the work of Szabó etĀ al. (2007), and lower precision data from Gaia DR3 for about 800 stars for which phase data are included in the database directly (Ripepi etĀ al., 2023).
Under the adiabatic assumption, one would expect radial pulsation to have a offset where maximum brightness corresponds to maximum compression, i.e, to zero pulsation velocity on the descent from RV maximum. In reality, the phase offset is closer to zero () for pulsating stars, because ionization removes heat energy during the compression phase, thus introducing a phase lag (Castor, 1968; Szabó et al., 2007). For Betelgeuse, we find a relatively large phase offset for the FM, although its amplitude dropped considerably after the Great Dimming, which makes the fit more uncertain. The phase offset for the O1 mode, which is currently dominating the light variations in Betelgeuse, agrees with that of the shorter-period pulsating stars.
To characterize the LSP beyond the specific context of Betelgeuse, we calculated the phase offsets for the LSP stars published by Nicholls etĀ al. (2009), where OGLE-III I-band photometry was available. We also calculated phase offsets for a subset of stars that were proposed to be LSP candidates by Gaia Collaboration etĀ al. (2023a), based on the recent Gaia FPR (Focused Product Release). Despite the fact that the Gaia data are generally sparse for these stars, we managed to process nearly a third (400 stars) of the full sample. The distributions of the two sets in FigureĀ 3 agree well, with most stars saddling the value. We also found a few targets with offsets close to .
Turning now to the binarity model of the LSP, the configuration proposed by SoszyÅski etĀ al. (2021) postulates that minimum light would occur when the companion passes in front of the red giant, corresponding to the point at which the RV variation of the primary goes through zero on the descending branch. This would occur at or offset. However, as FigureĀ 3 shows, we observe the opposite not only in Betelgeuse, but in the majority of the LSP stars.
4 Proposed LSP Hypotheses
In light of the observations discussed above, we now review and reevaluate proposed hypotheses for LSPs, discussing the validity of each as it applies to Betelgeuse.
4.1 Hypothesis 1: LSP is actually the FM
Recently, Saio etĀ al. (2023) proposed that the 2100-day periodicity observed in Betelgeuse is in fact Betelgeuseās fundamental pressure mode (FM) rather than an LSPāa claim that is in tension with a number of other recent investigations (Joyce etĀ al. 2020; Dupree etĀ al. 2022; NeuhƤuser etĀ al. 2022; Granzer etĀ al. 2022; MacLeod etĀ al. 2023; MolnĆ”r etĀ al. 2023; see also Wheeler & Chatzopoulos 2023 for a recent review). However, classifying the 400-d periodicity as the first overtone (O1) and the 2100-d periodicity as the FM cleanly resolves the question of the origin of the LSP in Betelgeuse by asserting that there simply isnāt one. This possibility is worth investigating.
Joyce etĀ al. (2020) provided evidence that the 2100-day periodicity cannot be the fundamental mode on the basis that a periodicity of this length is not activated by the -mechanism in either linear oscillation calculations performed with GYRE (Townsend & Teitler, 2013; Townsend etĀ al., 2018) or 1D hydrodynamic calculations performed with MESA (Paxton etĀ al., 2011, 2013, 2015, 2018, 2019; Jermyn etĀ al., 2023). Likewise, MacLeod etĀ al. (2023) demonstrate that since the Great Dimming, Betelgeuseās fundamental mode oscillation was perturbed into a combination of overtones. As in Joyce etĀ al. (2020), MacLeod etĀ al. (2023) identify the 400-d periodicity as the FM by performing adiabatic and non-adiabatic calculations with GYRE, exploring a larger range of spherical degrees than in Joyce etĀ al. (2020) (see their Table 1). Similarly, Jadlovský etĀ al. (2024) examine the behavior of shockwaves in the aftermath of the dimming, coming to the same conclusion as MacLeod etĀ al. (2023) that the dominant pulsation mode had transitioned from the FM to the first overtone.
In general, semiregular variable stars can exchange pulsational energy between -modes, so an exchange between the FM and the O1 is possible, but an exchange between the FM and LSP is not. In particular, energy exchange between modes happens in semiregular stars for period ratios close to a 2:1 resonant ratio, which is the ratio of the 400-d and 200-d periods of Betelgeuse (Buchler etĀ al., 2004). The fact that there is mode switch from the 400-day periodicity to a 200-d periodicity being observed right after the Great Dimming (also MacLeod etĀ al. 2023), rather than a mode switch at the 2100-day periodicity, is further evidence that the 400-d periodicity is the FM (as is the fact that 2100:400 is a 5:1 rather than 2:1 ratio).
Another critical issue with the 2100d-as-FM hypothesis is the stellar radius required to sustain an acoustic mode with a period of 2100 days. As noted in Saio etĀ al. (2023), a radius of roughly is implied by models that treat Betelgeuseās -day periodicity as the FM. Observational constraints on Betelgeuseās radius provided by Haubois etĀ al. (2009), MontargĆØs etĀ al. (2014), and Kervella etĀ al. (2018a) are in tension with these calculations, even at the longest distance measurements, as discussed in detail in MolnĆ”r etĀ al. (2023). Measurements that find angular diameters larger than 50 mas (corresponding to radii ) were done at longer wavelength and thus sample the MOLsphere (the close molecular layers) above the photosphere (OāGorman etĀ al., 2017; Cannon etĀ al., 2023).
Still other studies undermine the 2100d-as-FM hypothesis on strictly evolutionary grounds. For example, in an investigation focused on constraining Betelgeuseās present-day evolutionary stage, NeuhƤuser etĀ al. (2022) argue that Betelgeuse recently crossed the Hertzsprung gap, supported by its historically documented color change over the past several thousand years. A recent traversing of the Hertzsprung gap places Betelgeuse in the early stages of core helium burning, which is incompatible with the core-carbon-burning evolutionary phase implied by a 2100-day fundamental mode (Saio etĀ al., 2023).
Given the range of arguments against a 2100-d fundamental mode period from many orthogonal perspectives, we reject the claim that the pulsation mode widely classified as Betelgeuseās LSP is actually the starās FM. We proceed henceforth referring to the -d periodicity as the FM and the -d periodicity as the LSP in Betelgeuse. While the case of Betelgeuse alone is enough to debunk the 2100d-as-FM hypothesis as a universal explanation for LSPs, the existence of Sequence D on the PL diagram (see Figure 1) does so more conclusively.
4.2 Hypothesis 2: Giant convective cells
In light of the evidence that LSPs are not simply mis-identified FMs, as discussed above, Stothers & Leung (1971) proposed the hypothesis that giant convective cells can explain the LSP phenomenon, generalized by Stothers (2010) to extend to all LSP pulsators. Broadly, while the acoustic FM timescale is limited by and thus predicts periods too short to explain the LSP for reasonable stellar radii and , a hypothetical periodic signal might be mediated by the overturn timescale of giant convective cells,
where is the plume size and is the characteristic velocity of convection. In the limit of low with a maximum plume size approaching , this timescale can indeed be longer than the FM.
This picture has order-of-magnitude agreement in the luminosities and radial velocity amplitudes (Stothers, 2010). Following Mixing Length Theory (MLT, Böhm-Vitense 1958; see recent review by Joyce & Tayar 2023), the local characteristic length scale of convection can be expressed as where is the pressure scale height and for RSGs (e.g. Chun et al. 2018; Goldberg et al. 2022). In MLT, the can then be estimated (Kippenhahn et al., 2013) from the superadiabaticity , the local gravity , and thermodynamic properties
where is a geometric factor encoding plume geometry (typically following Henyey etĀ al. 1965 and others), is determined by the equation of state (EOS), and can be related to the flux carried by convection () and the mixing length by
and where is the specific heat at constant pressure.
In the extended convective envelopes of RGB stars, AGB stars, and RSGs, indeed . In such a star, large in the interior yields large-scale plumes spanning a sizeable fraction of the stellar radius, as recovered both by simulations (e.g. Freytag etĀ al., 2002; Dorch, 2004; Brun & Palacios, 2009; Plez & Chiavassa, 2013; Chiavassa etĀ al., 2009, 2010b, 2010a, 2011a, 2011b, 2017, 2024; Goldberg etĀ al., 2022; Antoni & Quataert, 2022) and inferred from observations (e.g. MontargĆØs etĀ al., 2017; Chiavassa etĀ al., 2018a, b; Kravchenko etĀ al., 2018, 2019, 2020, 2021; Zhang etĀ al., 2024). For RSGs like Betelgeuse, MLT predicts convective velocities of km/s for RSG stars (see discussions in Joyce etĀ al., 2020; Goldberg etĀ al., 2022), somewhat consistent with the required convective velocity to give rise to an overturn time of period in a star of radius
However, while there is order-of-magnitude agreement with the observed LSP, tensions arise. Large-scale convection is more or less ubiquitous in cool, luminous giants, and as such, we would expect all such stars to exhibit similar LSP behavior if giant convective cells were the cause. In reality, only 30% do (Nicholls etĀ al., 2009; SoszyÅski etĀ al., 2021). Moreover, convective motions do not remain coherent forever, even in the case of a giant convective cell, whereas LSPs are generally steady in nature. Takayama (2023), for example, found that of LSP candidates in the LMC and SMC, only of those candidates showed any evidence for a changing LSP duration over the observing term. In the convective picture, over time significant stochasticity is expected to be introduced into both the lightcurves and observed velocities, and eventually the stochastic variability should wipe out a periodic signal from convection alone. In fact, in the hotter OB star regime, variability from subsurface convection has been proposed to better explain stochastic low-frequency variability (Schultz etĀ al., 2020; Cantiello etĀ al., 2021), and in RSGs, similar stochastic variability from granulation has been recovered from LMC observations (e.g. Zhang etĀ al., 2024).
Giant convective cell models also make no prediction for RV modulation on the timescale of LSP. The convective velocity is a characteristic velocity of convective transport and not simply a bulk fluid velocity, as plumes move upwards and downwards and laterally at the surface (Stein & Nordlund, 1989, 1998). In fact, the convective overturn time is often considered to be a decoherence time. In 3D simulations of RSG envelopes Goldberg etĀ al. (2022) found decoherence times on the order of 300-500 days. Additionally, Ma etĀ al. (2024) show that convective plumes in an RSG-like convective envelopes could mimic a 5 km/s signal inferred to be surface rotation (see, e.g. Kervella etĀ al., 2018b). However, they find that the coherence time and dipole orientation of such a signal are significantly less than their 5 years of simulation time, with snapshots showing synthetic sin dipoles ranging from km/s to 10km/s at random orientations.
The MLT treatment which defines a single āconvective velocityā scale and thus a single āconvective overturnā timescale is, of course, a dramatic simplification of the true complexities of convection (see e.g. Joyce & Tayar 2023 §5, as well as the discussion in §5 of Goldberg etĀ al. 2022 in the case of RSGs, especially their Fig.Ā 19). Stars like Betelgeuse with large envelopes hosting vigorous convection are indeed more poorly approximated by MLT than solar-like stars by virtue of their significant deviation from solar conditions (see e.g. Joyce & Chaboyer 2018). Locally Super-Eddington luminosities due to the presence of strong opacity peaks lead to numerical as well as physical challenges for 1D models, which must be overcome through various āengineeringā techniques (see, e.g., §7.7 of Paxton etĀ al. 2013 and §7.2 of Jermyn etĀ al. 2023; and discussions and references therein). Moreover, the convective motions observed in 3D Radiation-Hydrodynamics simulations of near-surface convection in RSG envelopes other luminous massive stars do not even resemble motions typical of MLT-like convection out to the surface (Schultz etĀ al., 2020, 2022; Goldberg etĀ al., 2022); above some radius () the correlations between the radial fluid velocity and the density, entropy, and opacity invert ā cold, dense, high-opacity material which would sink in MLT-like convection is correlated with positive ā even at high optical depth .111 While there are significant caveats with MLT in the context of defining a single persistent convective overturn time, it is unlikely that the simplifying assumptions of MLT discussed above would affect results in this work in any other way. As noted by Joyce etĀ al. (2020), the primary impact of changing in models of 10-25 was to shift in temperature the location of the red giant branch, leading to a factor-of-ten increase in the uncertainty of the effective temperature constraint as compared to the error bars reported observationally by Levesque & Massey (2020). While uncertainty in the optimal value of the mixing length could certainly impact results through ās impact on the synthetic seismic frequency spectra against which Betelgeuseās pulsation spectrum was compared, this uncertainty was fully accounted for in Joyce etĀ al. (2020), whose parameters we use elsewhere in our analysis.
At the very least, any modulated RV signal caused by giant convective cell turnover is likely to be far more complex than the RV modulation seen in Betelgeuse or the Gaia LSP variables, and an LSP driven by large-scale turbulent motions would not persist for the hundred years over which Betelgeuse has been observed.
4.3 Hypothesis 3: Mode-interactions (beats)
Interaction between close-by modes can create long beating periods that could match the LSPs. However, beating modulates the pulsation amplitudes and hence the lightcurve envelopes, not in the average brightness. Furthermore, if the LSP period were a beating period between modes in Betelgeuse, then it would correspond to a physical signal close to the pulsation frequencies. This signal would be offset by about 19% and 10% relative to the FM or the O1 frequencies, instead being of an independent signal at .
Another aspect of mode interaction concerns interactions between the LSP and the other pulsation modes. If two (or more) strong pulsation modes are present in the star, they will both influence the envelope structure in which the other modes have to travel. These interactions would cause combination frequencies to appear in the frequency spectrum of the star, as seen in double-mode Cepheids and RR Lyrae stars (see, e.g. Gruberbauer etĀ al., 2007). Such combination frequencies with the LSP would be very similar to the beating frequencies mentioned above (e.g., at ). But such signals have not been reported for Betelgeuse, which suggests that the LSP is not the product of a beating phenomenon.
4.4 Hypothesis 4: Rotation
As noted in Wood etĀ al. (2004a), if stars exhibiting LSPs are ārotating prolate spheroids,ā meaning they are distorted in such a way that the polar radius is larger than the equatorial radius, they would produce the velocity curves characteristic of LSPs. However, the rotational velocities that would be required for this explanation are far too high compared to observations: the apparent 5-15 km/s surface velocity of Betelgeuse corresponds to a 30-year periodicity, whereas Betelgeuseās LSP has a 5.5 year periodicity. Further, LSPs fall on a period-luminosity sequence, and it is not clear why this would happen if (solid body) rotation were the explanation.
4.4.1 Differential Rotation
On the main sequence, the cores of high-mass stars have been predicted by some models to rotate about 100x as fast as their surfaces (see, e.g., Heger etĀ al. 2000, 2005), though more recent studies have shown that these models over-predicted core rotation by a factor of 10. In one such study, Pedersen (2022) found that measurements indicate that SPB stars with masses of 5-10 rotate close to rigidly. Conversely, Hermes etĀ al. (2017) do find evidence of a connection between high stellar birth mass and rapid core rotation via asteroseismology of white dwarfs, although again for masses below that of Betelgeuse. There are currently no measurements of core rotation in stars above , so extrapolation to Betelgeuse should be taken with caution.
Further, the relationship between envelope and core rotation along the main sequence is not necessarily preserved once the star enters the core helium-burning phase. Mosser etĀ al. (2024) find that, for low-mass red giants, the evolution of the core rotation rate in core-helium-burning stars scales with the inverse square of the stellar radius, in good agreement with Tayar etĀ al. (2019). Although Betelgeuse is a massive star, and the core rotation rates of red supergiants have not been probed directly, asteroseismic relations for low-mass, core helium-burning red giants are informative for this case. The model described in Mosser etĀ al. (2024) is that of a rigid, dense helium core surrounded by a radiative layer with smooth rotation profile, all encased by a rigid convective envelope: a picture that describes Betelgeuse reasonably well. Putting the Joyce etĀ al. (2020) lower mass bound () and radius () for Betelgeuse into Mosser etĀ al. (2024)ās equation 3 yields a core rotation rate of approximately 400 days. This is much shorter than the LSP, so modulation by internal rotational is ruled out.
Though Lovekin etĀ al. (2009) note that extreme cases of differential rotation will cause slight shifts in acoustic mode frequencies, the extent of such a shift would be nowhere near the factor of decrease in frequency required to make the FM mimic the LSP. Further, Jermyn etĀ al. (2020) find that there is very little radial differential rotation in stars with convective envelopes (in both short and long period eclipsing binaries) anyway.
4.5 Hypothesis 5: Magnetism
Various authors have explored the possibility that a process analogous to the magnetic solar cycleāe.g. star spots, convectively-driven magnetic dynamoācould generate LSPs (e.g. Soker & Clayton 1999 in AGBs and Wood etĀ al. 2004b in RSGs), but several problems arise with this theory (Olivier & Wood, 2003). Of particular importance is the fact that spots on a rotating surface are not capable of explaining lightcurve variations across the appropriate range of colors (i.e. the observed chromatic behavior, or chromaticity), as discussed in Wood etĀ al. (2004b) and more recently in Takayama etĀ al. (2015).
Another issue is that magnetic cycles in stars like Betelgeuse are expected to be much longer than 11 years. Although there is no simple means of scaling the solar magnetic cycle to red supergiants, and dynamo cycles have not been observationally confirmed in stars like Betelgeuse, we know that the period of the magnetic solar cycle goes as
where is the stellar radius and is the magnetic diffusivity. The magnetic diffusivity is a function of the scale of the convective motions in the envelope and the starās conductivity. Every term in this relation contributes to increasing the period of the magnetic cycle in Betelgeuse relative to that of the Sun. Given that the period of the LSP is roughly years, a magnetic cycle similar to the Sunās is not a viable explanation for the LSP in a large, evolved star with a deep convective envelope.
Some magnetic phenomena affecting high-mass stars are worth considering. The interplay between differential rotation and magnetic fields has been studied primarily in OB stars (Maeder & Meynet, 2011, 2012), which was Betelgeuseās spectral type when it was on the main sequence. (Keszthelyi etĀ al., 2019) find that fossil fields in massive stars can induce mass-loss quenching and magnetic braking; however, episodic mass loss is periodic and should happen at the frequency of the FM. The role of fossil fields would be to suppress the amount of mass dispersed per episode, not to alter the frequency of mass-loss episodes. Fossil fields therefore do not cause the LSP. Finally, while shorter-period, dynamo-generated magnetic fields can exist in the sub-surface convection zones in hot, massive stars (Cantiello & Braithwaite, 2011), Betelgeuse is not hot enough to have an iron opacity peak near the surface.
4.6 Hypothesis 6: Non-radial Pulsations
Any pulsation-related explanation for the LSP must involve a driving mechanism that produces pulsations below the frequency of the FM (see Section 4.1), which means it cannot be acoustically driven. Such drivers may be intrinsic or extrinsic to the star. Here we address the possibility that the LSP is caused by an intrinsic source of non-radial pulsation.
4.6.1 Strange Modes
Strange modes are unstable modes with potentially high growth rates that occur in extremely non-adiabatic environments dominated by radiative energy transport (Saio, 2009). They are caused by a large difference in phase between the density and pressure maximum in the oscillating partial ionization zone. Strange modes can be found in the radial spectra of many luminous stars, including Cepheids, RR Lyrae, LBVs, and other high-amplitude pulsators (Buchler etĀ al., 1997). Their presence in luminous red giants is discussed in detail in Wood & Olivier (2014).
Strange modes with LSP-like periods have been found in red giant models, arising due to the interaction of stellar oscillations with convective energy transport (Wood, 2000). However, such modes are highly damped, and models predict an exponential decay timescale of less than 100 days upon perturbation (Wood etĀ al., 2004b). They are therefore highly unlikely to be observed in real stars.
More recently, Saio etĀ al. (2015) argued that such oscillatory convective modes can, in fact, explain LSPs in real red giants. However, to generate modes of the correct period to reproduce Sequence D (see Figure 1), Saio etĀ al. (2015) engineered highly non-adiabatic stellar models dominated by radiative transport in their outermost layers, which do not accurately reflect the physics of red supergiants. Three-dimensional simulations show that in fully convective RSG envelopes, the convective velocities penetrate above the surface. Large turbulent motions well outside stellar photosphere form coherent plumes, meaning that while radiation carries much of the flux, the convective motions persist far above the photosphere (see, e.g. Chiavassa etĀ al., 2009, 2011b, 2024; Goldberg etĀ al., 2022; Ma etĀ al., 2024). In fact, at any radius defined as the āphotosphere,ā there is even some fluid that is optically thick and retaining its entropy such that its behavior is convective even though radiation is carrying flux elsewhere along in the near-surface layers (see discussion in SectionĀ 4.4 of Goldberg etĀ al. 2022).
Saio etĀ al. (2015) encoded the surface-radiation-dominated physics in their one-dimensional models (computed with MESA) by using a low, very sub-solar value for the convective efficiency () in the MLT formalism, and it is this choice that enables both the large surface radiative zones and large radii of their models. In reality, convection is more efficient in a red supergiantās envelope than it is in the Sunās (e.g. Chun etĀ al., 2018), with typical values of inferred from observed HR positions of nearby RSGs. Further, use of produces models with inappropriately low effective temperatures (Levesque & Massey, 2020).
Given the inability of models to reproduce simultaneously the correct RSG temperatures, convective envelope physics consistent with RSGs, LSP-appropriate strange mode frequencies, and strange mode lifetimes long enough to be observable, we reject strange modes as the explanation for the LSP of Betelgeuse.
4.6.2 g-modes
Because they can have longer periods than the radial fundamental mode, -mode pulsations were first proposed as an explanation for the LSP by Wood etĀ al. (1999a). However, the structure of red supergiant envelopes once again undermines the viability of this explanation. As noted by Takayama etĀ al. (2015) and discussed earlier in Wood etĀ al. (2004b), -modes must develop in a radiative zone. At best, LSP stars have a very thin outer radiative layer above their large convective envelopes. The size of this region strongly limits the amplitude of -modes that could develop there, and any such modes would be too small to explain the large amplitudes of both the brightness and radial velocity variations that characterize LSPs.
Inverting the argument, any -mode that would be observationally detectable as the LSP in Betelgeuse implies the existence of a quiet, āwell-behaved,ā and large-enough radiative zone near the surface of Betelgeuse in which this mode is propagating. Neither stellar evolution calculations nor models of large convective cells in red supergiants suggest this is possible.
To comment briefly upon -modes in the stellar core, we note that -modes are evanescent (i.e. exhibiting amplitude decay over distance) throughout the convective envelopeāwhich is, once again, enormous in Betelgeuseāand would therefore not be detectable at the stellar surface. We are thus left with the fact that -modes can account for LSP periods but not LSP amplitudes, no matter their origin in RSGs.
4.7 Hypothesis 7: -mechanism in the dust shell
The final remaining pulsational scenario consistent with the LSP timescale posits a radial pulsation from a mechanism driven outside the optical stellar photosphere, dragging both the photosphere and the external material. Winters etĀ al. (1994) and Hoefner etĀ al. (1995) argue that opacity changes in the dust shell surrounding the star could cause oscillations via a ādust-ā mechanism, whereby the strong temperature and density sensitivity of the sublimation state (and thus opacity) of circumstellar dust could induce radial pulsations in the dust shell(s) (see also discussions by Wood etĀ al., 2004b).
From a period-mean-density relation with generic period ,
where is the average density and the second proportionality assumes the mass scale is dominated by the stellar material and not the dust.222In order to achieve a modulated RV signal, near stellar-photosphere material would need to take part in the oscillation, so the stellar mass is expected to set the mean density unless there is very appreciable mass within the dust shell. For an LSP:FM period ratio of , assuming the FM is an oscillation at the (optical or IR) stellar photosphere, this would mean that the āouterā radius mediating the RV signal would be times the stellar radius, or for a radius of (Joyce etĀ al., 2020). This is consistent with placement inside the region probed by radio observations indicating the presence of significant circumstellar dust from stellar radii (Altenhoff etĀ al., 1979; Lim etĀ al., 1998; Harper etĀ al., 2001; Smith etĀ al., 2009; OāGorman etĀ al., 2015).
A dust--like mechanism is also appealing because it simultaneously predicts the chromatic behavior of the LSPs and could explain IR lightcurve dips apparently modulated by increases and decreases in dust activity in front of the star (SoszyÅski, 2007; SoszyÅski etĀ al., 2021; Percy & Shenoy, 2023). Such a pulsation, if it is indeed dragging the whole star, would also explain the typical LSP LC-RV phase offset, with the LSP luminosity minimum (indicating more dust) occuring shortly after the maximum in the modulated stellar RV (indicated by low-but-increasing photospheric RV).
Ultimately, the ability of such a pulsation out at 3 to adequately modulate the RV at the level of km/s at the stellar surface is the primary point of tension for the dust- theory. As pointed out by Wood etĀ al. (2004a, b), if dust is being created and ejected via a dust- mechanism, this phenomena would occur primarily outside the stellar photosphere and it is unclear how the deeper photospheric layers could be dragged enough to cause the RV amplitude modulations seen in LSPs, and this argument extends also to Betelgeuse. Moreover, changes in and the optical stellar photosphere would be accompanied by large changes in . We note that this was not seen even in the Great Dimmingās extreme dust ejection near LSP minimum (MontargĆØs etĀ al., 2021), during which Betelgeuseās cooled slightly but remained at K (Levesque & Massey, 2020). This low variation in is consistent with other LSP variables, which show little modulation in (Wood etĀ al., 2004b). Finally, though Betelgeuse does exist within a surrounding dusty CSM, it should be noted that some LSP stars show little evidence for large circumstellar dust shells to begin with (Hinkle etĀ al., 2002; Olivier & Wood, 2003). While some kind of externally-sourced pulsation remains capable in theory of reconciling the RV phase and the LSP, there is no clear way to do this without large corresponding changes in .
4.8 Hypothesis 8: Modulation by a companion
In this section, we overview the ways in which a companion could induce an LSP (see, e.g. Wood etĀ al., 1999a, 2004b; Retter, 2005; SoszyÅski, 2007; SoszyÅski etĀ al., 2021; Percy & Shenoy, 2023, and discussions and references therein). This family of solutions is the generally favored mechanism for producing LSPs, as this class of scenarios naturally explains the long timescale and steady period as the orbital time of the companion. As discussed in SectionĀ 2, the existence of a P-L relation for LSPs is not an intrinsic problem, as orbital architecture is sensitive to the stellar mass and radius. A remaining point of tension is the relatively high fraction of long-period variables with LSPs and RV modulations compared to main-sequence stars (Nicholls etĀ al., 2009), which can be addressed by appealing to mass growth of the orbiting body (e.g. Livio & Soker 1984; see discussion in §5 of SoszyÅski etĀ al. 2021). At least regarding RSGs, the observed binary fraction is very high for massive stars (Sana etĀ al., 2012; de Mink etĀ al., 2013), and this fraction does not include undetected lower-mass companions (we return to this in SectionĀ 5.2).
4.8.1 Hypothesis 8a: Tidally-Induced Oscillations
Tidally-induced oscillations are another form of non-radial oscillation (see Section 4.6) that emerge in highly non-adiabatic conditions (see, e.g., Bunting etĀ al. 2019; Sun etĀ al. 2023). In the case of a companion inducing modes on Betelgeuse via tidal forces, it is the companionās orbital timescale that could modulate the RV variations, but an oscillation of the star or system that determines the luminosity fluctuations. If we consider tidally-driven modes induced on Betelgeuse by a companion, these modes are still acoustic -modes acting over the surface of Betelgeuse and are therefore limited by its radius (). Considering the case of a severely tidally distorted Betelgeuse, because at the surface, to distort the star substantially in a Roche potential would require a very short orbital period on the order of the FM. As such, tidally-driven modes in the star itself cannot have significantly lower frequencies than the fundamental mode.
If, on the other hand, we consider a scenario in which the potential of the orbital systemāBetelgeuse and its companionādrives a pulsational mode that acts over the radius of the whole system (rather than Betelgeuse alone), it is unclear what would be pulsating. If the oscillating medium is a steady-state dust shell encompassing both bodies, then we encounter difficulties providing a physical justification for a pulsation in such a low-density environment (see also SectionĀ 4.7). There remains the possibility that such an oscillation might mediate the properties of the dust, thereby modulating the luminosity and chromatic behavior. If the orbit provides RV modulation in 1:1 resonance with an induced oscillation period, this might explain why there could be a sizeable RV modulation with relatively steady . In general, however, this scenario would not prefer any particular phase relationship between the LSP and the RV, since a pulsating dust shell modulating the luminosity would not know about the observerās direction. In either case, tidally-induced oscillations do not provide a clean explanation for the LSP.
4.8.2 Hypothesis 8b: Occultation
The current preferred hypothesis for LSPs posits that a low-mass companion is dragging a cloud of dust and gas along the observerās line of sight, temporarily blocking the view of the star and causing the luminosity minimum (see, e.g. Wood etĀ al., 1999b, 2004b; SoszyÅski, 2007; SoszyÅski & Udalski, 2014; SoszyÅski etĀ al., 2021; Percy & Shenoy, 2023).
This theory is appealing as it
(1) naturally explains a timescale longer than the FM,
(2) allows for RV modulation on the LSP timescale
(3) without substantial changes in , and
(4) explains dust-like chromatic behavior.
As discussed in detail by SoszyÅski etĀ al. (2021), a compelling piece of evidence for dust modulation on an orbital timescale as an explanation for the LSP phenomenon comes from inspecting the shape of the LSP lightcurves. These vary in the characteristics of their luminosity fluctuations, from narrower dips commensurate with small orbiting dust blobs, through deep periodic signals consistent with much larger or more complex orbiting dusty bodies. As discussed in SectionĀ 2, there is no significant tension between a companion hypothesis and the existence of a P-L relation; in fact, this succinctly explains why Sequence D in our FigureĀ 1 and Fig.Ā 2 of SoszyÅski etĀ al. (2021) extends from the sequence of lower-luminosity eclipsing binaries (sequence E).
A diversity of LSP amplitudes would be expected for different inclination angles and companion/dust cloud properties, and this model can explain cases where there is no detectable RV signal by invoking a very high mass ratio. However, the occultation hypothesis predicts that when RV modulation is present on the LSP timescale, one should infer the orbiting companion to be somewhere along the observerās line of sight, i.e., the starās should be near-average and accelerating towards the observer near luminosity minimum, at a phase offset of or (see FigureĀ 3).
Unfortunately, this prediction does not hold for Betelgeuse with its LSP-RV phase offset of . This is almost out of phase with what would be expected for dust gravitationally bound to an orbiting body or at the L1 Lagrange point and still out of phase with dust sitting the L4 or L5 Lagrange points. The occultation hypothesis is also in tension with the RV phases for the majority of the Gaia sample shown in FigureĀ 3 and discussed in SectionĀ 3. Thus while the orbiting dusty body hypothesis can explain the RV amplitude and LC period in Betelgeuse, it does not explain their phase offset.
4.8.3 Hypothesis 8c: Circumstellar Modulation Near Companion
| Hypothesis | Section | Timescale | Properties | Low variation | Dust-like | LC-RV | Persistence |
| of RV | in | chromaticity | offset | ||||
| Misidentified FM | 4.1 | ā+ | X | ? | - | ? | ā+ |
| Giant convective cells | 4.2 | ā | ? | ā | - | - | X |
| Mode interactions | 4.3 | X | X | - | - | - | X |
| Rotation | 4.4 | X | ā | ? | ? | ā | ā+ |
| Magnetism | 4.5 | X | - | X | X | ? | ? |
| Non-radial pulsation | 4.6 | ā+ | ā | ? | ? | ? | X |
| Dust- | 4.7 | ā+ | ā+ | X | ā+ | ā+ | ā |
| Binarity: tidal | 4.8.1 | X | ā+ | ? | ? | ? | ā+ |
| Binarity: occultation | 4.8.2 | ā+ | ā+ | ā+ | ā+ | X | ā+ |
| Binarity: dust modulation | 4.8.3 | ā+ | ā+ | ā+ | ā+ | ā+ | ā+ |
The tension between the dusty companion hypothesis and the RV/lightcurve phase offset requires reevaluating the broader physical picture. From FigureĀ 2, the phase difference between the lightcurve and RV modulation favors a scenario in which the photosphere begins to move away from the observer near the time of LSP luminosity minimum. For this to be caused by a companion, the companion must be 180 degrees out of phase with the dust that induces the brightness minimum of the LSP. Thus we propose a new scenario: a companion whose orbital period sets the timescale of the LSP and RV fluctuations, and whose presence modulates, removes, eliminates, or changes the optical depth of any potentially dusty material in its vicinity, rather than dragging a trailing cloud of dusty material. There are multiple ways in which this might occur, including, for example, dynamically perturbing the potential of the dusty medium or through irradiation.
For this scenario to be feasible, there must be material for the companion to interact with, and the timescale over which the circumstellar matter replenishes and enters into proximity with the companion must be shorter than the orbital period. Both of these conditions appear to be satisfied in Betelgeuse. As supported by observational measurements of material extending from the MOLsphere to many tens of stellar radii beyond, it is well-established that dusty material exists outside the photosphere (Smith etĀ al., 2009; Kervella etĀ al., 2009, 2011; MontargĆØs etĀ al., 2014; OāGorman etĀ al., 2015; Dharmawardena etĀ al., 2018; Haubois etĀ al., 2019).
The connection between binarity, convection, pulsation, dust condensation, and mass loss has been explored by various groups in both simulations and observations (see, e.g., Höfner & Olofsson 2018 for a recent review in the context of AGB mass loss). Both 1D and 3D simulations of evolved cool giant stars exhibit dust-driven winds forming at a few times the stellar radius, seeded by (asymmetric) material which gets shocked and levitated by the pulsating stellar surface (see, e.g., Freytag & Höfner 2008; Freytag et al. 2017; Freytag & Höfner 2023; Liljegren et al. 2016, 2017, 2018; Bladh et al. 2019a, b; Siess et al. 2022; Steinwandel et al. 2022; Wiegert et al. 2024). The outflows in these works exhibit a wide range of circumstellar densities and mass-loss rates depending properties near the stellar surface, such as the pressure scale height, temperature, pulsation period, and degree of asymmetry (as well as composition; see, e.g. Karakas & Lattanzio 2014; Lattanzio & Karakas 2016).
Likewise, observations indicate a correspondence between the onset of dusty winds and different pulsation sequences (e.g. McDonald & Trabucchi, 2019; Yu etĀ al., 2021). LSPs have long been associated with dusty activity, as Wood & Nicholls (2009) found increased mid-IR emission with no difference in color for RGB stars with LSPs compared to those without, interpreted to indicate the presence of dust in a clumpy or disk-like configuration. It has also been shown in hydrodynamic simulations of AGB star winds that a companion may modify the circumstellar matter formed by dusty outflows (Maes etĀ al., 2021), with differing geometry depending on the properties of the outflow and companion.
Two important timescales may set the launching of periodic or stochastic outflows which might turn into dusty fodder: pulsations and convection. The pulsation timescale (FM or O1) is necessarily shorter than the LSP, and the coupling of near-sonic, large-scale convection in a pulsating envelope has been hypothesized to drive stellar winds Yoon & Cantiello (2010). Similarly, the convective motions themselves may lead to steepening shocks in the atmosphere. This may stratify material beyond the stellar photosphere, well past the dust formation radius, with variations changing over the convective turnover time Fuller & Tsuna (2024). A timescale for this process can be estimated by considering motions in the deep interior, following equation B4 of Ma etĀ al. (2024),
where is the density at the optical photosphere g/cm3, or surface timescales much shorter than that (See discussion in Appendix B of Ma etĀ al. 2024). It is also known that for particularly extreme dust ejection events, such as the Great Dimming (MontargĆØs etĀ al., 2021; Dupree etĀ al., 2022; Drevon etĀ al., 2024) a major dust cloud forms and passes on timescales shorter than the LSP (see extensive discussion in MacLeod etĀ al., 2023).
The modification we are presenting to the dust-modulating-companion scenario originally discussed in, e.g., Wood etĀ al. (1999a); SoszyÅski (2007); SoszyÅski & Udalski (2014); SoszyÅski etĀ al. (2021) is appealing for the four reasons mentioned in SectionĀ 4.8.2 (timescale, RV modulation, low variance in , and chromaticity), with the additional benefit that it explains the phase offset in Betelgeuse, as well as many of the Gaia LSP variables (FigureĀ 3).
A dust-modulating companion also leaves open the possibility of secondary eclipses to be seen in the IR, which represent strong evidence for the binarity scenario in other cases (SoszyÅski etĀ al., 2021). The primary difference is that the companion would be 180 degrees out of phase with the bulk of the circumstellar dust rather than dragging the dust along as in the occultation scenario. This phase difference would potentially but not necessarily yield differences in the (highly-uncertain) expected dust geometry, depending in part on whether the dust modulation is dynamical or radiative in nature.
We note that elimination or dispersal of dust near an orbiting body need not be the explanation for all LSP variables. Many LSP variables do show steady amplitudes and periodic narrow lightcurve dips (SoszyÅski etĀ al., 2021), and at least a few show RV modulation phases consistent with luminosity drops when the companion would be in front of the star. It is likely that the mass, luminosity, and of the companion, as well as the replenishment timescale of material provided by the star, may determine whether the system drags and condenses dust along its orbit or creates an under-density in its surroundings. We leave it to future work to determine the exact conditions under which each orbitally-modulated LSP/RV phase scenario (occultation, dynamical dispersal, accretion, irradiation, etc.) could arise.
We thus conclude that circumstellar dust modulation by a companion is the most plausible explanation for the LSP in Betelgeuse. Within this framework, we can use the observational data available to estimate the properties of a companion consistent with the RV and RV-LC phase difference. This companion is hereafter referred to as 333āBā stands for Betelbuddy. The properties (SectionĀ 5) and detectability (SectionĀ 5.2) of are discussed subsequently.
4.9 Summary of Hypotheses
TableĀ 2 provides a quick-look reference for the explanatory power and limitations of each theory discussed in this section.
5 Companion-as-LSP in Betelgeuse
In Section 4, we argued that the most plausible explanation for the LSP in Betelgeuse is a low-mass companion () that impacts the dust in its vicinity, corresponding to a brightness increase when the companion is in view, a novel revision of the scenario described by SoszyÅski etĀ al. (2021). FigureĀ 4 gives a schematic diagram of the Ori system proposed in SectionĀ 4.8.3 over one cycle of the LSP, and illustrates the approximate location of at various points in the phase-folded lightcurve/RV cycles from FigureĀ 2: (A) increasing RV at LSP minimum ( behind Betelgeuse), (B) maximum RV at LSP median ( on the limb moving into view), (C) decreasing RV at LSP maximum ( in view), and (D) minimum RV at LSP median ( on the limb moving out of view). In the remainder of this section, we constrain the orbital properties of the posited , and explore the implications of such a companion.
5.1 Properties of Companion
From the RV data we estimate the mass of via the binary mass function:
where is the velocity amplitude and ād is the LSP period from the RV observations of Granzer etĀ al. (2022), and is the present-day mass estimate of from Joyce etĀ al. (2020). For simplicity, we assume a circular orbit. This assumption is further justified given the expected effects of dust reducing eccentricity, investigated primarily in the context of planets in dusty disk environments by, e.g. Li etĀ al. (2019), Coleman etĀ al. (2022). This gives us a lower mass limit of . This value is an order of magnitude larger than the mass estimates calculated for LSPs in RGBs with RV constraints, which are typically in the or less range, i.e., closer to brown dwarfs than to dwarf stars (Nicholls etĀ al., 2009). In order to cause substantial modulation over the course of an orbit, we assume that a relatively low inclination relative to our line of sight (or a high inclination relative to the plane perpendicular to our line of sight, as defined conventionally for binary stars and exoplanets) is needed for any interaction with the circumstellar dust to be observable from our vantage point. Therefore, we estimate that the mass of is unlikely to exceed , or at most with a vertically extended cloud.
Using this mass constraint, we calculate the orbital separation assuming a Keplerian orbit:
which recovers an orbital separation of assuming the present-day mass of Betelgeuse estimated by Joyce etĀ al. (2020). Using their stellar radius of yields an orbital distance of . This is well outside of the CO2-H2O MOLsphere described by MontargĆØs etĀ al. (2021), which extends to . It also places the L1 Lagrange point at , or . In this orbital configuration, the surface of Betelgeuse is not expected to deform significantly, as the Roche radius for this orbit is , and the equipotential corresponding to is nearly spherical. From the calculated separation, we can estimate a transit probability of , and a corresponding minimal inclination of for a grazing transit of (where an inclination of means an edge-on orbit). Therefore, even with an extended dust feature that can reach higher inclinations, our assumption of low inclination relative to the observerās line of sight is justified.
It remains unclear from these orbital parameters alone whether is able to modulate dust. Any irradiation effects will depend on the companionās age and outgoing radiation field and the dust properties, which dynamics alone cannot recover (see further discussion in SectionĀ 5.2). From dynamical arguments, one can calculate the Hill sphere of the companion star, which approximates its sphere of gravitational influence,
This gives within the likely ranges for ās mass, or of the radius of Betelgeuse. This is at least a plausible angular cross section to subtend over which one might see modulation.
5.2 Observability of Companion
Even at its highest plausible massāsay āit would be nearly impossible to detect close to such a luminous and intrinsically variable star (the average brightness of Betelgeuse hovers near ) except in some exotic and improbable scenarios (see SectionĀ 5.2.1). Even adopting an optimistic 10 for , the luminosity contrast would still be , likely undetectable in integrated light. This challenge is exacerbated by the fact that Betelgeuseās intrinsic variability from convection and pulsations entails mag fluctuations in the V-band (see our Fig.2), with intrinsic fluctuations on the order of more than , at least 1000 times the intrinsic brightness of .
Spectral differencing at different phases would be another possible avenue for detection, but again, subtle changes in the of Betelgeuse and changes on the FM timescale would overwhelm variations in spectra unless the companion were substantially hotter than Betelgeuse. At its hottest, a 2 star might be spectral class F, corresponding to a of K. A companion of this temperature would only just barely be visible on the Wien tail in the UV (a 7000K blackbody outshines a brighter, 3600K blackbody only above Hz, or below Ć ), and realistically, this is a drastic overestimate of the temperature of .
5.2.1 Exotic Scenarios
While the companionās mass and proximity to a massive star might lead one to speculate that it may be a neutron star or other exotic compact body, as far as the authors of this study are aware, no such detections have been made (Posson-Brown etĀ al., 2006; Kashyap etĀ al., 2020). The current lack of X-ray evidence notwithstanding, there are a few reasons not to entirely ignore a neutron-star-as-companion scenario. First, it is more common than not that high-mass stars such as Betelgeuse are born as binary or even triple systems (Sana etĀ al., 2012; de Mink etĀ al., 2013; Toonen etĀ al., 2018, 2020, 2022; Offner etĀ al., 2023), many of which will produce runaway stars when the most massive star in the system explodes (e.g. Eldridge etĀ al., 2011; Zapartas etĀ al., 2021). Second, the anomalous proper motion of Betelgeuse (e.g. Harper etĀ al., 2008) still requires explanation, and a neutron star companion could provide one. In that scenario, the progenitor to a neutron star born at the same time as Betelgeuse must have been more massive than Betelgeuse in order to die earlier in a Supernova explosion. The past SN explosion of Betelgeuseās companion would have then been energetic enough to provide the necessary ākickā that sent Betelgeuse along its present observed trajectory, with the binary widening but without disrupting the system (Renzo etĀ al., 2019). Future X-ray observations scheduled at favorable epochs might rule this out conclusively. Such an experiment could also help characterize the presence of a possible corona around a non-compact companion, which might yield insight into the exact mechanism for interaction with any circumstellar dust.
Using an approximate age for Betelgeuse of Myr, and assuming Betelgeuse and its companion are co-natal, it is possible that could be a low-mass, pre-main sequence star hosting an accretion or proto-planetary disk. While this could be bright enough to detect, given typical disk lifetimes between years (Fiorellino etĀ al., 2023), the probability of catching the system in this particular configuration is exceedingly low. The authors are likewise unaware of any existing observational evidence of this scenario.
5.2.2 Observability Under Normal Conditions
For spectral differencing or photometry, the closeness of the orbit can be exploited, since the probability of the companion being eclipsed from our vantage point is . Therefore, an eclipse in the UV passbands would be potentially detectable. Future space-based UV spectroscopic missions such as UVEX (Kulkarni etĀ al., 2021) would be well-positioned to investigate this. However, one still would have to contend with the length of an eclipse. As for the orbital parameters and a near-circular orbit, the approximate time for to cross the disk of Betelgeuse is days, i.e., longer than an observing season, and comparable to the FM and O1 pulsation periods.
Another possibility is to inspect the shape of evolving asymmetries in sub-milimeter observations of the dusty region around Betelgeuse using an instrument like ALMA. There is indeed substantial asymmetry in the dust observed near the radius of the predicted orbital separation of (e.g. Richards etĀ al., 2013; OāGorman etĀ al., 2017, 2020; Kervella etĀ al., 2018b).
As the sensitivity and achievable luminosity contrast of instruments with powerful coronagraphs and extreme adaptive optics systems designed for planet-hunting and other purposes improve, we are optimistic that in the future such a companion will be detectable, and in the meanwhile we emphasize this system will make for an interesting target for creative observers.
For future targeted observations, the RV phase can be used to predict where will be relative to the star over the next LSP cycle and beyond. Directly observing a transit of is impossible given Betelgeuseās intrinsic periodic and stochastic variability from convection and pulsations on shorter timescales than the transit. However, the times when is near the limb could in theory reveal its presence and stellar properties if observed at sufficiently high contrast and spatial resolution. Given the LSP period of 2170 days, the following are the upcoming epochs where would be near maximum separation:
ā¢
Min RV: JD2460651 06/12/2024
ā¢
Max RV: JD2461736 26/11/2027
ā¢
Min RV: JD2462821 15/11/2030
ā¢
Max RV: JD2463906 04/11/2033
with syzygy halfway in between.
| Parameter | Value | Reference |
| Radius of Betelgeuse | Joyce etĀ al. (2020) | |
| Mass of Betelgeuse | Joyce etĀ al. (2020) | |
| Radius of Betelgeuse | Joyce etĀ al. (2020) | |
| Orbital Period | 21695.3d | Granzer et al. (2022); Jadlovský et al. (2023) |
| of | This work | |
| Orbital Separation | This work | |
| D/M/Y of next RV min | 06/12/2024 | This work |
| D/M/Y of next RV max | 26/11/2027 | This work |
| D/M/Y of second RV min | 15/11/2030 | This work |
| D/M/Y of second RV max | 04/11/2033 | This work |
In the absence of instrumentation breakthroughs enabling extremely high luminosity-contrast direct imaging for stars as bright as Betelgeuse (), we believe the best chance for observability or falsification is with repeated targeted radio-interferometric observations continuously throughout the LSP cycle. These studies should look specifically for patterns in the dust modulation and asymmetry correlated with ās expected orbital separation.
A similar observing strategy was undertaken by OāGorman etĀ al. (2015), who took VLA measurements at 0.7, 1.3, 2.0, 3.5, and 6.1cm at multiple different phase points along the LSP over a 4-year cycle (in addition to single-epoch observations by Lim etĀ al. (1998) 4 years prior). Their Fig.Ā 1 does indicate some time-dependent modulation of an asymmetric circumstellar envelope visible in the 0.7cm interferometric data around mas444At maximum separation, will be at mas for a stellar optical photosphere diameter of 43 mas), but it is far from conclusive evidence, as an 8-year time span entails measurements during multiple different LSP cycles (orbits of ). While we do expect intrinsic diversity in dust structures on a replenishment timescale commensurate with the FM as the FM interacts with the large-scale convection, would cause a systematic modulation in the circumstellar dust (the proposed mechanism for the LSP itself).
If an apparent under-density of dust migrates around the star on the LSP cycle, and if that matches the points in the LSP cycle where should be on the limb versus transiting Betelgeuse versus in opposition, this would lend strong evidence to the existence of as proposed. Such a finding would also further constrain the systemās orbital plane and inclination, the companionās stellar properties, and the mechanism by which interacts with the dust.
6 Conclusions
In this study, we have demonstrated that the most likely explanation for the prominent 2170-d periodicity observed in Betelgeuse is a low-mass, binary companion, dubbed , that modulates dust. The proposed orbital picture leads to reduced obfuscation (higher apparent stellar ) when is in transit.
In Section 2, we reviewed the variability phenomenon known as Long Secondary Periods (LSPs) present in roughly 30% of all long-period variable stars and in Betelgeuse specifically. In Section 3, we compiled and presented new data (Figure 2) comparing Betelgeuseās lightcurve and RV cycle side-by-side and characterized the phase offset between them. We showed that the phase offset present not only in Betelgeuse, but in the majority of Gaia LSP stars, contradicts the physical picture set by the prevailing theory for binarity as the driver for LSPs proposed by SoszyÅski etĀ al. (2021).
In light of the tension introduced by the RV-lightcurve phase offset, we extensively investigated the explanatory power of every hypothesis proposed for LSPs in the literature as applied to Betelgeuse. Section 4 covered possibilities ranging from large-scale convection, magnetism, mode interactions, non-radial pulsations, and exotic physics, but only converged on one viable explanation: Betelgeuseās LSP is caused by a low-mass companion, , in a dusty environment.
In Section 5, we calculated physical properties of the companion and the Orionis orbital system as constrained by existing observations. A summary of these parameters is provided in TableĀ 3. We discussed why has not been detected before, commented on exotic physical scenarios that would be theoretically observable, discussed practical limitations on the companionās detectability, and described the advances in method, technique and instrumentation needed to detect it in the future.
While our work was in the refereeing process, MacLeod etĀ al. (2024) independently found astrometric evidence that a companion scenario can explain Betelgeuseās LSP, focusing their efforts on dissecting Betelgeuseās long baseline RV observations. They recovered similar orbital parameters and a slightly lower mass () due to a smaller inferred RV amplitude.
The confirmation of would have far-reaching implications for Betelgeuse and its evolutionary future. Betelgeuse is expected to explode as a Type II-Plateau Supernova, and when it does, it could appear nearly as bright as the quarter-moon (Goldberg etĀ al., 2020a). There is substantial uncertainty in the expected early supernova luminosity evolution due to the unknown status of Betelgeuseās pulsation phase at the time of explosion (Goldberg etĀ al., 2020b; Hsu etĀ al., 2024) and the properties of the near-surface material (e.g. Morozova etĀ al., 2017, 2018; Hiramatsu etĀ al., 2021; Moriya etĀ al., 2011, 2018; Kozyreva etĀ al., 2022; Moriya etĀ al., 2023; Jacobson-GalĆ”n etĀ al., 2024).
The presence of dense circumstellar material is of reinvigorated interest after the recent Supernova 2023ixf, the closest core-collapse event of the decade, which showed strong signatures of early interaction in an asymmetric circumstellar medium (e.g. Hiramatsu etĀ al., 2023; Jacobson-GalĆ”n etĀ al., 2023; Hosseinzadeh etĀ al., 2023; Smith etĀ al., 2023; Teja etĀ al., 2023; Zimmerman etĀ al., 2024; Li etĀ al., 2024; Martinez etĀ al., 2024; Singh etĀ al., 2024; Moriya & Singh, 2024; Bostroem etĀ al., 2024) around a pulsating Red Supergiant progenitor (e.g. Soraisam etĀ al., 2023; Jencson etĀ al., 2023; Kilpatrick etĀ al., 2023; Xiang etĀ al., 2024). Moreover, supernova remnants are now being observed with unprecedented resolution in the era of JWST (e.g. recent work by Temim etĀ al., 2024; Milisavljevic etĀ al., 2024), and the proposed observations of Betelgeuseās surrounding environment, whether or not they support the presence of , will be valuable in constraining the expected level and geometry of circumstellar asymmetries that may imprint on the remnant morphology (Polin etĀ al., 2022; Mandal etĀ al., 2023, 2024). If the existence of is confirmed and a progenitor identified, one could also predict its own future evolution in the context of Betelgeuseās eventual explosion. Such simulations have recently been carried out in the case of Type Ia SNe with a surviving companion (Bauer etĀ al., 2019; Bhat etĀ al., 2024; Wong etĀ al., 2024).
Observations of Betelgeuse have identified photometric variations on a 30 year period consistent with a rotational km/s (Joyce etĀ al., 2020), which for an inclination angle of 20ā recovers an equatorial velocity of 15Ā km/s (see Dupree etĀ al. 1987; Gilliland & Dupree 1996; Uitenbroek etĀ al. 1998; Kervella etĀ al. 2009, 2018b, and review byWheeler & Chatzopoulos 2023). This has led to speculation that Betelgeuse may be a merger product (e.g. Wheeler etĀ al., 2017; Chatzopoulos etĀ al., 2020; Shiber etĀ al., 2024). This hypothesis was invoked to account for fast envelope rotation which would not survive the spin-down as Betelgeuse ascended up the RSG branch in a single-star evolutionary scenario. Although the large-scale convective envelope itself might be the source of this RV signal (Ma etĀ al., 2024), if the apparent RV dipole signal persists at a consistent inclination or at higher ALMA resolution, another possible explanation might be angular momentum transferred from the orbit of , which could be informed by further constraining the stellar mass and properties of .
In this picture, it is not purely a coincidence that the Great Dimming of 2020 (e.g. Dupree etĀ al., 2020, 2022; MontargĆØs etĀ al., 2021; Drevon etĀ al., 2024) occured near LSP Minimum. As the opening angle of such an ejection is expected to be large if caused by a large pressure perturbation from convection in the interior (MacLeod etĀ al., 2023), a transiting companion would likely disrupt its path. However, a companion on the other side of Betelgeuseās potential would likely not interact. One may speculate that dusty mass loss could be funneled through the orbitās L3 lagrange point, though dynamical modeling including gas-dust interactions are necessary to quantiatively evaluate this claim.
Appendix A Data tables
In this section we present the following data in numerical format. TableĀ 4 lists the brightnesses, distances and periods for the Galactic supergiant stars from Fig.Ā 1. TableĀ 5 lists the phase offset values we calculated for the stars presented by Nicholls etĀ al. (2009), whereas TablesĀ 6 and 7 contains the phase offset for the subset of Gaia LSP candidates that we include in Fig.Ā 3.
| ID | d | P1 | eP1 | P2 | eP2 | P3 | eP3 | Ref | |||
| (mag) | (pc) | (mag) | (d) | (d) | (d) | (d) | (d) | (d) | |||
| SS And | 0.971 | 159 | 17 | 1 | |||||||
| NO Aur | 3.729 | 362 | 11 | 38.4 | 0.3 | 1 | |||||
| VY CMa | 0.291 | 1600 | 190 | 1 | |||||||
| RT Car | 1.500 | 201 | 25 | 448 | 146 | 1 | |||||
| CL Car | 1.539 | 490 | 100 | 229 | 14 | 2600 | 1000 | 1 | |||
| EV Car | 0.788 | 276 | 26 | 820 | 230 | 1 | |||||
| IX Car | 1.884 | 408 | 50 | 4400 | 2000 | 1 | |||||
| TZ Cas | 1.939 | 3100 | 1000 | 1 | |||||||
| PZ Cas | 0.781 | 850 | 150 | 3195 | 800 | 1 | |||||
| ST Cep | 1.644 | 3300 | 1000 | 1 | |||||||
| Cep | -1.620 | 860 | 50 | 4400 | 1060 | 3 | |||||
| T Cet | -0.808 | 298 | 3 | 161 | 3 | 2 | |||||
| RW Cyg | 0.640 | 580 | 80 | 1 | |||||||
| AZ Cyg | 1.288 | 495 | 40 | 3350 | 1100 | 1 | |||||
| BC Cyg | 0.299 | 720 | 40 | 1 | |||||||
| TV Gem | 0.947 | 426 | 45 | 2550 | 680 | 1 | |||||
| W Gem | 1.850 | 353 | 24 | 1 | |||||||
| BU Gem | 0.806 | 2450 | 750 | 1 | |||||||
| Her | -3.511 | 124 | 5 | 500 | 50 | 1480 | 200 | 3 | |||
| W Ind | 3.108 | 193 | 15 | 1 | |||||||
| Y Lyn | -0.688 | 133 | 3 | 1240 | 50 | 1 | |||||
| XY Lyr | -0.213 | 122 | 1 | ||||||||
| Ori | -4.378 | 410 | 32 | 200 | 20 | 2100 | 200 | 5 | |||
| S Per | 1.123 | 813 | 60 | 4 | |||||||
| T Per | 2.581 | 2500 | 460 | 1 | |||||||
| W Per | 1.568 | 500 | 40 | 2900 | 300 | 1 | |||||
| RS Per | 1.562 | 4200 | 1500 | 1 | |||||||
| SU Per | 1.455 | 430 | 70 | 3050 | 1200 | 1 | |||||
| XX Per | 1.972 | 3150 | 1000 | 1 | |||||||
| BU Per | 2.194 | 381 | 30 | 3600 | 1000 | 1 | |||||
| FZ Per | 2.482 | 368 | 13 | 1 | |||||||
| VX Sgr | -0.122 | 754 | 56 | 4 | |||||||
| AH Sco | 0.415 | 738 | 78 | 1 | |||||||
| Sco | -4.100 | 1650 | 640 | 3 | |||||||
| CE Tau | -0.913 | 1300 | 100 | 1 | |||||||
| W Tri | 1.091 | 107 | 6 | 590 | 170 | 1 |
| OGLE ID | (d) | e | |
| LMC-LPV-55909 | 783.7 | -1.07 | 0.21 |
| LMC-LPV-55739 | 562.1 | -1.04 | 0.38 |
| LMC-LPV-55861 | 1007 | -1.95 | 0.11 |
| LMC-LPV-55565 | 810.2 | -2.34 | 0.16 |
| LMC-LPV-57329 | 1032 | -2.22 | 0.48 |
| LMC-LPV-56928 | 800 | -1.21 | 0.15 |
| LMC-LPV-56063 | 540 | -3.16 | 0.43 |
| LMC-LPV-57356 | 434.8 | -1.74 | 0.17 |
| LMC-LPV-57073 | 470.5 | -1.72 | 0.21 |
| LMC-LPV-58350 | 611.6 | -2.17 | 0.13 |
| LMC-LPV-58399 | 818.7 | 3.90 | 0.13 |
| LMC-LPV-58174 | 640.9 | 3.96 | 0.11 |
| LMC-LPV-59753 | 1048 | 4.36 | 0.29 |
| LMC-LPV-59287 | 257.7 | -3.02 | 0.46 |
| LMC-LPV-60196 | 1139 | -1.93 | 0.35 |
| LMC-LPV-61999 | 404.3 | -3.07 | 0.21 |
| LMC-LPV-60643 | 535.8 | -1.02 | 0.20 |
| LMC-LPV-61577 | 720.5 | -1.75 | 0.15 |
| LMC-LPV-62722 | 715.8 | -2.48 | 0.17 |
| LMC-LPV-63825 | 849.5 | -2.86 | 0.27 |
| Gaia DR3 ID | (d) | e | Gaia DR3 ID | (d) | e | ||
| 5366556864650293504 | 467.8 | -3.56 | 1.98 | 4518782615952131200 | 413.4 | -1.98 | 0.52 |
| 5245519394861788416 | 424.3 | -3.38 | 1.95 | 4591864301979656960 | 384.0 | -1.95 | 0.83 |
| 2030155380681518080 | 415.3 | -3.09 | 1.70 | 1958845802441119360 | 445.3 | -1.70 | 0.24 |
| 4877073701212505984 | 467.9 | -3.07 | 1.62 | 5823469916709143808 | 433.2 | -1.62 | 0.93 |
| 1009476541884049152 | 436.2 | -3.00 | 1.53 | 5404420815320630528 | 561.6 | -1.53 | 0.56 |
| 416537855493120896 | 592.8 | -2.54 | 1.48 | 1863461622559594624 | 472.3 | -1.48 | 0.73 |
| 5592948576556271360 | 567.5 | -2.54 | 1.47 | 2030200671149815424 | 442.6 | -1.47 | 0.37 |
| 5487656658716777984 | 523.9 | -2.52 | 1.43 | 5823227714910793856 | 368.1 | -1.43 | 0.27 |
| 4242144837569110656 | 560.2 | -2.52 | 1.36 | 5819332488801340928 | 381.8 | -1.36 | 0.28 |
| 5193827058256042752 | 780.7 | -2.47 | 1.35 | 718681245623682176 | 446.5 | -1.35 | 0.35 |
| 2940744561181222784 | 510.5 | -2.46 | 1.31 | 5592924490378357632 | 427.0 | -1.31 | 0.46 |
| 5393712362310861824 | 378.9 | -2.46 | 1.21 | 2022032910328304128 | 488.3 | -1.21 | 0.15 |
| 5365837715327919872 | 294.4 | -2.44 | 1.20 | 2226627353963447552 | 507.6 | -1.20 | 0.73 |
| 5193010395993757952 | 585.3 | -2.42 | 1.18 | 6036139001931775360 | 403.9 | -1.18 | 0.39 |
| 1863379262264189568 | 339.0 | -2.42 | 1.18 | 413371051550885248 | 436.6 | -1.18 | 0.27 |
| 4325121337971096704 | 325.6 | -2.42 | 1.11 | 5842150894131928576 | 367.5 | -1.11 | 0.23 |
| 6648842337535201152 | 304.8 | -2.42 | 1.04 | 5615714067889118336 | 605.4 | -1.04 | 0.37 |
| 5783448380630645760 | 462.0 | -2.41 | 1.04 | 419317730125131904 | 434.7 | -1.04 | 0.25 |
| 525948917547462272 | 770.1 | -2.39 | 1.03 | 5464229059521244928 | 289.8 | -1.03 | 0.41 |
| 3417134843227320192 | 814.3 | -2.38 | 0.96 | 4325728161012091904 | 453.3 | -0.96 | 0.44 |
| 5870612332427790464 | 328.5 | -2.37 | 0.95 | 2223228905258736128 | 680.9 | -0.95 | 0.94 |
| 5793428029206907136 | 466.0 | -2.36 | 0.95 | 5836237308379605248 | 387.0 | -0.95 | 0.32 |
| 4216847961220342400 | 353.1 | -2.35 | 0.90 | 2057374688680289408 | 685.6 | -0.90 | 0.91 |
| 2162840771749435904 | 701.0 | -2.34 | 0.87 | 563548504362254464 | 547.5 | -0.87 | 0.32 |
| 6065282485339565440 | 694.1 | -2.33 | 0.87 | 2111189048343657856 | 498.7 | -0.87 | 0.19 |
| 4333627052547566464 | 638.3 | -2.31 | 0.86 | 6076773820652573312 | 397.6 | -0.86 | 0.36 |
| 6066743976821731200 | 354.2 | -2.31 | 0.82 | 5779249135267388160 | 462.7 | -0.82 | 0.26 |
| 4373808155229106048 | 415.7 | -2.29 | 0.81 | 487598707882598912 | 333.9 | -0.81 | 0.82 |
| 5814559886782561152 | 483.4 | -2.28 | 0.81 | 2033040091219687424 | 495.0 | -0.81 | 0.17 |
| 1562772473976590976 | 345.3 | -2.27 | 0.78 | 2082877414211373056 | 514.7 | -0.78 | 0.58 |
| 4167947800049312128 | 379.9 | -2.27 | 0.75 | 4994896061572855680 | 242.8 | -0.75 | 0.29 |
| 4382027799216609280 | 366.1 | -2.26 | 0.73 | 5350758909803889152 | 584.9 | -0.73 | 0.38 |
| 5985892282645097728 | 335.5 | -2.24 | 0.73 | 1384487655667958528 | 778.5 | -0.73 | 0.57 |
| 5994696415836802304 | 687.5 | -2.21 | 0.69 | 2038497375395267328 | 489.4 | -0.69 | 0.28 |
| 5614629674553165568 | 594.9 | -2.20 | 0.68 | 4755154010768073088 | 355.0 | -0.68 | 0.62 |
| 2928002527094437504 | 678.3 | -2.20 | 0.67 | 5357213966450101760 | 749.1 | -0.67 | 0.29 |
| 5693199740370024576 | 363.9 | -2.19 | 0.66 | 2004024043737361280 | 618.2 | -0.66 | 1.03 |
| 6011162495794543104 | 304.5 | -2.18 | 0.65 | 4660125660364324864 | 750.0 | -0.65 | 0.23 |
| 5241758687171977216 | 704.5 | -2.17 | 0.64 | 5717734865611449984 | 412.1 | -0.64 | 1.16 |
| 6702743863564684800 | 393.9 | -2.16 | 0.63 | 5730565341664785792 | 455.5 | -0.63 | 0.31 |
| 2924804941112026240 | 525.0 | -2.14 | 0.63 | 426958575728796032 | 536.2 | -0.63 | 0.14 |
| 4487425953063740800 | 683.4 | -2.10 | 0.38 | 5646561553800726784 | 265.4 | -0.38 | 0.29 |
| 5914192369148073216 | 460.6 | -2.09 | 0.33 | 4490750562689619584 | 273.9 | -0.33 | 0.35 |
| 5846845430839141632 | 482.8 | -2.07 | 0.29 | 4599913964043089408 | 380.3 | -0.29 | 0.19 |
| 2192933679121399808 | 674.3 | -2.02 | 0.17 | 5856898196963666816 | 692.4 | -0.17 | 0.29 |
| 5970189980950373248 | 482.1 | -2.02 | 0.15 | 2016488970841656320 | 510.8 | -0.15 | 0.29 |
| Gaia DR3 ID | (d) | e | |
| 2021971062805508224 | 595.7 | -0.09 | 0.30 |
| 515032034094986112 | 663.0 | -0.06 | 0.25 |
| 2026398010175881856 | 490.4 | 0.29 | 0.61 |
| 5243641635146938112 | 350.2 | 0.54 | 0.30 |
| 6076305364288485888 | 310.9 | 0.57 | 0.05 |
| 2231448029552393472 | 394.8 | -4.85 | 0.15 |
| 4688988660900813952 | 370.2 | -4.83 | 0.58 |
| 6641986087963535872 | 350.5 | -3.71 | 0.27 |
| 5348049300824404864 | 476.4 | -3.58 | 0.44 |
| 513141801805888896 | 712.0 | -3.21 | 0.30 |
| 2198382038425517696 | 720.6 | -3.21 | 0.30 |
| 5834193689811972480 | 466.2 | -3.16 | 0.27 |
| 1975711383110161664 | 464.2 | -3.10 | 0.23 |
| 5328193048536364672 | 384.6 | -3.08 | 0.40 |
| 6649474625438931840 | 289.4 | -3.00 | 0.65 |
| 2231049353507901568 | 547.2 | -2.52 | 0.27 |
| 5328530461163971968 | 514.8 | -2.41 | 0.82 |
| 5829828839147242240 | 435.6 | -2.41 | 0.26 |
| 1632704127922550272 | 447.1 | -2.36 | 0.44 |
| 2883038617538998144 | 518.9 | -2.35 | 0.26 |
| 5646198474439605760 | 539.1 | -2.34 | 0.15 |
| 6723996873576976768 | 324.7 | -2.30 | 0.25 |
| 1496998382733052928 | 321.9 | -2.28 | 0.30 |
| 6076713656747162624 | 623.0 | -2.27 | 0.45 |
| 2061975320222333696 | 544.2 | -2.19 | 0.65 |
| 5693299903313069312 | 318.1 | -2.11 | 0.28 |
| 4601169090926461184 | 384.2 | -1.93 | 1.42 |
| 2162696838801402624 | 507.8 | -1.57 | 0.21 |
| 5998083014713314944 | 398.2 | -1.56 | 0.44 |
| 385255959408720000 | 474.3 | -1.31 | 0.33 |
| 2030233136772978304 | 444.6 | -1.14 | 0.55 |
| 2895408329510160384 | 564.8 | -1.09 | 0.73 |
| 4617724574943762560 | 501.1 | -1.01 | 1.05 |
| 2045973161114595584 | 310.2 | -0.96 | 0.24 |
| 506098605197562624 | 570.6 | -0.94 | 0.24 |
| 6091162171552477440 | 406.6 | -0.91 | 0.20 |
| 4569516728144529664 | 406.0 | -0.89 | 0.35 |
| 4487550915135532288 | 411.8 | -0.81 | 0.19 |
| 4613334534251566848 | 534.7 | -0.80 | 0.28 |
| 513591398982018944 | 735.0 | -0.75 | 0.30 |
| 2014407045572567424 | 354.5 | -0.68 | 0.46 |
| 186324619130226048 | 575.0 | -0.67 | 1.66 |
| 505958902802490752 | 589.8 | -0.67 | 0.60 |
| 5638830642723891328 | 655.7 | -0.65 | 0.12 |